Fermi National Laboratory


Accelerator Update for February 9, 2005 - February 11, 2005

Wednesday February 9
The day shift began with Operations monitoring store 3972, with the Main Injector (MI) off due to switch-gear problems, with the Antiproton Source (Pbar) stack at 40.3E10 and being monitored, and with MiniBooNE taking beam.

The Neutron Therapy Facility (NTF) requested beam at 11:49 AM for patient treatment.

At 11:50 PM, a Booster RF station (BRF4) tripped off. It reset okay.

Thursday February 10
The midnight shift began with Operations monitoring store 3972, with MI off, and with MiniBooNE taking beam.

Operations terminated store 3972 at 5:30 AM.

Operations turned off the Booster beam at 6:56 AM. This shut off MiniBooNE beam.

Cryo system technicians began work on the Tevatron (TeV) sector E3 wet engine at 8:07 AM.

By 8:34 AM, operators had prepared Booster, MI, Pbar, Switchyard (SY), Meson, and the TeV for accesses.

Operations turned the Booster back on at 3:45 PM.

Linac experts began a study period at 4:42 PM.

NTF finished using beam at 6:11 PM.

Operations established beam to MiniBooNE at 9:57 PM.

Friday February 11
The midnight shift began with MiniBooNE taking beam.

At 3:52 AM, BRF15 tripped off. It reset okay.

Operations issued keys at 7:40 AM for an access into the TeV F sector. This access will inhibit MiniBooNE beam.

The Plan for Friday and the Weekend
The plan for today is to complete the MI work. With the understanding that many things must come together properly, we're hoping to have beam in MI early Saturday morning, with Pbar stacking after that, and then with Operations establishing a store later in the weekend.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/11/2005   email Fermilab